Please help me solve this equation:
The sum of the least and greatest of three consecutive integers is 72. What are the integers?

Answer by Cromlix(4381)

hi there,
Three consecutive integers
n, n + 1, and n + 2
The sum of the least and greatest
of three consecutive integers is 72.
n + n + 2 = 72
Collect like terms:
n + n = 72 - 2
2n = 70
n = 35
Three consecutive integers
35, 36 and 37
Hope this helps :-)